OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) Working together, two people can cut a large lawn in 5 hr. One person can do the job alone in 1 hr less than the other. How long would it take the faster person to do the job?

OpenStudy (legends):

Did i get it right las time @shoplovepc

OpenStudy (legends):

last*

OpenStudy (anonymous):

I didn't see an answer... I need a number lol

OpenStudy (holly00d1248):

3 and 2 you dont remember that

OpenStudy (anonymous):

Yeah it told me that was wrong :(

OpenStudy (legends):

thats wierd does it have any answer choices

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

type in answer

OpenStudy (legends):

wow:l

OpenStudy (anonymous):

Two people can cut one lawn in 5 hour; on their own, it takes approximately twice that, so 10 hours. 10 hours + 10 hours = 20 hours A + B = 20 A = B+1 Substitute the value of A (B+1) + B = 20 2B + 1 = 20 Subtract 2B = 19 B = 9.5 If B = 9.5 and A is an hour slower, A = 10.5 . The faster person can cut the lawn by himself in 9.5 hours.

OpenStudy (legends):

x = time of faster person 5/x + 5/(x+1) = 1 10x+5 = x^2+x x^2-9x-5 = 0 x = [9±sqrt(101)]/2 x = 9.525 hours

OpenStudy (legends):

lol all I see is a bunch of numbers that don't make sense

OpenStudy (anonymous):

lol that's ok... Can you help with one other problem?

OpenStudy (legends):

Sure

OpenStudy (anonymous):

Working together, two people can cut a lawn in 3 hr. One person can do the job alone in 1 hr less than the other. How long would it take the faster person to do the job?

OpenStudy (anonymous):

Basically the same but with 3 hours lol

OpenStudy (legends):

Oh ok I was about to tell you I think its the same lol.

OpenStudy (anonymous):

Wait nvm i got it! :) Thanks for all your help though

OpenStudy (legends):

no problem can you fan? :)

OpenStudy (anonymous):

Just did! :)

OpenStudy (legends):

Cool any other question just tag my name

OpenStudy (anonymous):

Okay! :)